- 1-[ : 13-1:-)-E -6 17]

determine whether the set vector is M2,2 is linearly independent or linearly dependent
Answer #1

o A: [3] wo - 4 -8 B = C - [ -3 17 5. CIA + C2 B + Ca C = 0 7 ([:]+013;]+91-8?). * 201 - 402 -8 (320 (i) -C2 313=0 (۱۱) (111